When a substance is synthesized under laboratory conditions, the probability of an


explosion in a separate experiment is p = 0.02. Determine the probability that in a series of 10


syntheses: 1) the explosion will occur three times; 2) the explosion will not occur

1
Expert's answer
2022-01-17T16:47:52-0500

"n=10 \\

p = 0.02 \\

q = 1 -0.02 = 0.98 \\

P(X=x)=C(n,x)p^x q^{n-x}"

1) the explosion will occur three times

"P(X=3) = \frac{10!}{3!(10-3)!} \times 0.02^3 \times 0.98^{10-3} = 0.000833"

2) the explosion will not occur

"P(X=0) = \frac{10!}{0!(10-0)!} \times 0.02^0 \times 0.98^{10-0} = 0.817072"